OpenStudy (anonymous):

Solve this system of linear equations. Seperate the x- and y- values with a comma. -x=62-18y 2x=38-18y

OpenStudy (anonymous):

Any clue?

OpenStudy (anonymous):

@ganeshie8 @mathstudent55

OpenStudy (anonymous):

\[-x=62-18y~~\iff~~x=18y-62\] Substitute into the second equation: \[2(18y-62)=38-18y\] Solve for \(y\).

OpenStudy (anonymous):

Ok